OpenStudy (anonymous):

Factor the trinomial completely. a^4-2a^2-35

OpenStudy (anonymous):

Do you know how to factor in general?

OpenStudy (anonymous):

i know you have to look for a number that when you add you get 2 and when you multiply you get 35

OpenStudy (anonymous):

but when i tried it out i got it wrong

OpenStudy (anonymous):

Any idea what those two numbers are?

OpenStudy (anonymous):

1 and 5?

OpenStudy (anonymous):

Hmm does the product of 1 and 5 get you close to 35 or even add up to two?

OpenStudy (anonymous):

If not, what are some multiples of 35 that we can assess?

OpenStudy (anonymous):

5 and 7 is 35 and the only number i can think that gets 2 is 1. or does it have to be the same number to get 35 and 2?

OpenStudy (anonymous):

No it doesn't have to be the same number. Bingo, you hit it :)

OpenStudy (anonymous):

5 and 7 is your answer

OpenStudy (anonymous):

Now put it in the equation to factor

OpenStudy (anonymous):

Remember it's negative two :)

OpenStudy (anonymous):

oh! so it would be (x+5) (x+7)?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Your first line should look like this a^4 +5a^2 - 7a^2 -35

OpenStudy (anonymous):

ahh i don't get it :/

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Ah dw :) I'll help you understand it

OpenStudy (anonymous):

Ready?

OpenStudy (anonymous):

yes and thank you :)

OpenStudy (anonymous):

Alright so you have that whole idea of factoring. How to factor? Simply find two numbers that add up to the middle and multiply to the last. Right? Alright, so the situation of factoring is not always the same. It's different. The answer may not look right to you, but it's right when you multiply it back up again.

OpenStudy (anonymous):

In our case, we have an equation with a degree of 4.

OpenStudy (anonymous):

How do we find it? Nothing changed. Same steps, different numbers.

OpenStudy (anonymous):

Just like opening a potato bag, Different shapes, still the same method of cooking it.

OpenStudy (anonymous):

Ready to move into the question?

OpenStudy (anonymous):

perfect. yes

OpenStudy (anonymous):

Alrighty :) So moving into the equation: a^4 - 2a^2 -35 We found two numbers, yeah? 7 and 5. 5 - 7 = -2 5*7 = 35 Perfect. All we gotta do now is put it into the equation.

OpenStudy (anonymous):

Btw for the subtraction, 7 is a negative. Hence, 5+(-7) = x 5 - 7 = x -2 = x

OpenStudy (anonymous):

You always add the numbers. The numbers however are not always positive.

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

oh ok yes got it

OpenStudy (anonymous):

now it's starting to make sense

OpenStudy (anonymous):

Alright that's what I wanted to do. :)

OpenStudy (anonymous):

Now that you have your two numbers, we insert them into the equation and factor :)

OpenStudy (anonymous):

Wanna try that? Show me step by step so you don't confuse yourself. Remember same steps, different numbers. Just follow the equation. Don't force an answer that appears orthodox.

OpenStudy (anonymous):

idk if i'm right.. um is it (a^2+5)(a^2-7)?

OpenStudy (anonymous):

I would kiss you congrats, but I can't. AWESOME MAN. You learned well :)

OpenStudy (anonymous):

You're right :)

OpenStudy (anonymous):

Remember what I taught you, keep up the good work.

OpenStudy (anonymous):

hahahaha! yay :) awesome thank you so much

OpenStudy (anonymous):

Ahaha anytime Ma'am. See you around :)

OpenStudy (anonymous):

see ya :)